business finance

At what annual interest rate, compounded annually, would

$

have to be invested for it to grow to

$

in

years?

The annual interest rate, compounded annually, at which

$

must be invested for it to grow to

$

in

years is

nothing%.

(Round to two decimal places.)
Answer #1

Solutions Given data fv = future value $1,934.16 PV = presont value =$ 490 R= Rate of Interest = ? n no. of years = 10 n Fv=

answered by: ANURANJAN SARSAM
Answer #2

SOLUTION :


Let the annual rate be r (in decimals) at which $480 invested today will grow to $1966.65 in 14 years by annual compounding.


So, 


FV = PV( 1+ r)^n

=> 1966.65 = 480(1 + r)^14

=> r = (1966.65/480)^(1/14) - 1

=> r = 0.1060 

=> r = 10.60 % 


The annual rate would be 10.60% . (ANSWER)


answered by: Tulsiram Garg
Answer #3

SOLUTION :


Let the annual rate be r (in decimals) at which $480 invested today will grow to $1966.65 in 14 years by annual compounding.


So, 


FV = PV( 1+ r)^n

=> 1966.65 = 480(1 + r)^14

=> r = (1966.65/480)^(1/14) - 1

=> r = 0.1060 

=> r = 10.60 % 


The annual rate would be 10.60% . (ANSWER)


answered by: Tulsiram Garg